OpenStudy (anonymous):

HELP PLEASE! (i give medals) Graph the system of inequalities on your paper. Please provide only one ordered pair (x, y) that will work for both system of inequalities. show all work y>-4x+1 y<2x+2

OpenStudy (goformit100):

y>-4x+1 y<2x+2

OpenStudy (anonymous):

-4x+1=2x+2 -6x=1 x=-1/6 y=5/3

OpenStudy (anonymous):

@JeanetteBaker can you explain the steps please?

OpenStudy (anonymous):

set the two equations equal and solve for x, then use that x value to solve for y in one of the equations, that is your x,y point